OpenStudy (anonymous):

Find the first five non-zero terms of power series representation centered at x=0 for the function below. f(x)=ln(6−x)

OpenStudy (amistre64):

it might be useful to know the power series for ln(1-x) to start with.

OpenStudy (anonymous):

Oh I got the answer ln(6 - x) = ln 6 - Σ(n = 0 to ∞) x^(n+1)/[6^(n+1) * (n+1)], which has R = 6

OpenStudy (amistre64):

thats fair

OpenStudy (anonymous):

I'm just confused on finding the first five terms?

OpenStudy (amistre64):

ln(6-x) the derivative is -1/(6-x) which can be long divided to produce the derivatives power series -1/6 -x/6^2 -------------- 6-x | -1 (-1+x/6) ------- -x/6 (-x/6+x^2/6^2) -------------- -x^2/6^2 i see a pattern \[\frac{-1}{6-x}=\sum_0\frac{-1}{6^{n-1}}x^n\] \[\int \frac{-1}{6-x}=\int\sum_0\frac{-1}{6^{n-1}}x^n\] \[ln(6-x)=\sum_0\frac{-1}{6^{n-1}(n+1)}x^{n+1}\]

OpenStudy (amistre64):

pull out the first 6 terms; n=0 to 5

OpenStudy (amistre64):

hmm, first term seems of course is ln(6-0) for the constant

OpenStudy (anonymous):

the second term is -x^2 / 2

OpenStudy (amistre64):

1/6^(n+1) slipped of the finger

OpenStudy (amistre64):

no \[ln(6-x)=ln(6)+\sum_0\frac{-1}{6^{n+1}(n+1)}x^{n+1}\] \[ln(6-x)=ln(6)+\frac{-1}{6^{0+1}(0+1)}x^{0+1}+\sum_1\frac{-1}{6^{n+1}(n+1)}x^{n+1}\] \[ln(6-x)=ln(6)-\frac{1}{6}x+\frac{-1}{6^{1+1}(1+1)}x^{1+1}+\sum_2\frac{-1}{6^{n+1}(n+1)}x^{n+1}\] etc ... ln(6) - x/6

OpenStudy (amistre64):

just pull out terms

OpenStudy (amistre64):

ln(6) - x/6 -x^2/2(6^2) - x^3/3(6^3) .... see a pattern?

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

@amistre64 when ur done i need u

OpenStudy (amistre64):

i believe im done, but please do not post for help i another persons question, its considered rude.